# UAV-VLPA*: A Vision-Language-Path-Action System for Optimal Route Generation on a Large Scales

### v1

@@ Abstract
The UAV-VLPA* (Visual-Language-Planning-and-Action) system represents a cutting-edge advancement in aerial robotics, designed to enhance communication and operational efficiency for unmanned aerial vehicles (UAVs). By integrating advanced planning capabilities, the system addresses the Traveling Salesman Problem (TSP) to optimize flight paths, reducing the total trajectory length by 18.5% compared to traditional methods. Additionally, the incorporation of the A* algorithm enables robust obstacle avoidance, ensuring safe and efficient navigation in complex environments. The system leverages satellite imagery processing combined with the Visual Language Model (VLM) and GPT’s natural language processing capabilities, allowing users to generate detailed flight plans through simple text commands. This seamless fusion of visual and linguistic analysis empowers precise decision-making and mission planning, making UAV-VLPA* a transformative tool for modern aerial operations. With its unmatched operational efficiency, navigational safety, and user-friendly functionality, UAV-VLPA* sets a new standard in autonomous aerial robotics, paving the way for future innovations in the field.
Index Terms:
VLA; VLM; LLM-agents; VLM-agents; UAV; Navigation; Drone; Path Planning.
@@ I Introduction
Figure 1: The pipeline of the UAV-VLPA system.
The field of aerial robotics has advanced significantly, with unmanned aerial vehicles (UAVs) becoming crucial in various applications, including surveillance, agriculture, and disaster management[1]. As UAV missions grow more complex, the need for effective path planning and autonomous mission generation has become increasingly important. Traditional UAV control methods typically rely on manual waypoint navigation or basic pre-programmed paths[2],[3], which are often labor-intensive and require specialized knowledge to operate effectively. However, as missions become more dynamic and environments more complex, these traditional methods are no longer sufficient, and more intelligent systems are needed for efficient and autonomous mission planning.
In recent years, Vision-Language-Action (VLA) and Vision-Language-Navigation (VLN) systems have emerged as an innovative solution, enabling UAVs to autonomously navigate from natural language instructions[4], [5], [6], [7]. These systems are based on transformer models that interpret natural language inputs and, in some cases, use satellite data to plan and execute missions[7]. Building upon the UAV-VLA framework[4], which generates actionable paths from linguistic instructions and satellite images, this paper extends the system by incorporating the Traveling Salesman Problem (TSP) into the mission generation process to ensure optimal path planning.
Leveraging the language request, we focus on introducing path planning for UAVs by integrating obstacle avoidance techniques with TSP for optimal route generation. The objective is to reduce mission completion time and minimize trajectory length while ensuring safe navigation through environments with potential obstacles. Using the outputs of the UAV-VLPA* system, we apply TSP to optimize the routes, taking into account environmental constraints and minimizing errors in path execution. This enables the UAV to generate more efficient paths that avoid obstacles while ensuring mission objectives are met.
The integration of satellite data into the UAV-VLPA* system significantly enhances its capabilities, as such data is widely accessible and open globally. By utilizing open satellite data alongside real-time environmental information, the system can improve the accuracy and adaptability of path planning and obstacle avoidance. This combination allows for more informed decision-making, enabling UAVs to navigate complex environments with greater precision and efficiency. The use of open satellite data not only enriches the system’s contextual understanding but also ensures cost-effective and scalable solutions for diverse aerial missions.
We use the benchmark UAV-VLPA-nano-30 [4] to evaluate the performance of our system, demonstrating that the integration of TSP with obstacle-aware path planning results in improved mission execution, outperforming human-level performance in large-scale UAV missions. Our contributions are as follows:
  • We introduce the UAV-VLPA* framework, combining Vision-Language based Path Planning with the Traveling Salesman Problem and Path Planning for optimal route generation.
  • We apply obstacle avoidance using A* algorithm in the mission generation process to ensure safe and efficient UAV navigation in complex environments.
  • We demonstrate that our integrated system, outperforms human-level path planning in large-scale UAV mission scenarios.
@@ II Related Work
The introduction of Vision Transformers (ViT) [8],[9] marked a significant advancement in the development of full-fledged models capable of processing and integrating multiple types of input and output, including text, images, video, and more. Building on this progress, the leading companies introduced models like ChatGPT-4 Omni[10] and DeepSeek R1 [11], which can reason across audio, vision, and text in real time, enabling seamless multimodal interactions. To address the problem of objects finding in robotics applications, Allen Institute of AI introduced model Molmo, that can point the requested objects on an image[12].
The usage of the transformer-based models allowed the extensive developing of the new methods, benchmarks, and datasets for Vision Language Navigation and Vision Language Action tasks. Firstly, the problem of Aerial Visual Language Navigation was proposed by Liu et al. [13], where they introduced the Aerial VLN method together with AerialVLN dataset. In [7] Fan et al. described the simulator and VLDN system, that can support the dialog with an operator during the flight. Lee et al.[5] presented an extended dataset with geographical meta information (streets, squares, boulevards, etc.). The introduction of dataset was paired with the new approach for goal predictor. Zhang et al. [14] took a pioneering step by building a universal environment for embodied intelligence in an open city. The agents there can perform both VLA and VLN tasks together online. Gao et al. [15] presented a method, where a map was provided as a matrix to the LLM model. In that work was introduced the Semantic Topo Metric Representation (STMR) approach, that allowed to feed the matrix map representation into the Large Language Model. In [16] Wang et al. presented the benchmark and simulator dubbed OpenUAV platform, which provides realistic environments, flight simulation, and comprehensive algorithmic support.
@@ III System Overview
Originally, the Visual Language Model (VLM) provided outputs by identifying point objects on an image in a sequential manner, from left to right and top to bottom, without considering obstacles or their spatial relationships. To address these limitations, we propose the following steps.
@@ III-A System Workflow
    Figure 2: The Waypoint Generation
@@
    Figure 3: The 2-Opt Local TSP Search
@@
    Figure 4: The A* algorithm
@@ III-C Path Planning and Obstacle Avoidance
To efficiently execute the A* algorithm in large-scale environments, the map is discretized into a uniform grid of equally sized cells, where each cell represents a node in the graph. The grid resolution is determined by balancing path fidelity—how accurately the path adheres to real-world terrain and obstacles—and computational efficiency—the number of nodes processed during the search. In this work, the discretization step is set to 5 pixels.
@@ IV Experimental Setup
The system was evaluated using the command: “Create a flight plan for the quadcopter to fly around each building at a height of 100 m, return to home, and land at the take-off point. Avoid lakes”. The experiment was conducted on a PC with an RTX 4090 graphics card (24GB VRAM) and Intel Core i9-13900K processor. Due to memory constraints, the quantized Molmo-7B-D BnB 4-bit model [23] was used.
@@ V Experimental Results
V-A Results with only TSP
To evaluate the impact of the Traveling Salesman Problem (TSP) on the UAV-VLA system, we applied TSP to the UAV-VLPA-nano-30 benchmark. The total trajectory length generated by the system was 51.34 km, which is significantly shorter than the trajectory length from the human-generated flight paths, which was 63.89 km, as reported in the UAV-VLA paper [4]. The manual flight plans used in the UAV-VLA paper are consistent with the results presented here, and only the TSP optimization was implemented, resulting in a reduction of the trajectory length from 77.74 km to 51.34 km.
The performance of the system was further evaluated using various error metrics, including KNN error, DTW RMSE, and Interpolation RMSE [4]. The results are summarized in Table I:
TABLE I: Comparison of RMSE Metrics for Different Methods on UAV-VLA with TSP
@@ V-A Results with only TSP
Mean 45.1586 259.066 354.463
Median 26.8865 235.483 308.907
Max 336.22 719.478 734.317
Additionally, we present two key plots to visualize the trajectory length comparison and the error analysis. Fig. 5 shows the comparison of trajectory lengths generated by the UAV-VLA system[4] with TSP and by an experienced operator, highlighting the efficiency gains when applying TSP. Fig. 6 illustrates the error distribution of the UAV-VLA system with TSP in comparison with ground truth, showing the performance of the system across different error metrics.
Figure 5: The comparison of the trajectory lengths made by UAV-VLA with TSP and by an experienced operator.
Figure 6: The error of UAV-VLA with TSP in comparison with ground truth.
The results demonstrate that integrating TSP with the UAV-VLA system significantly enhances the efficiency of UAV path planning, reducing trajectory lengths while maintaining high performance across error metrics.
V-B Results with the A* path planning algorithm
For this and next sections human-made path was changed due to additional requirement of avoiding the lakes on images. That’s why total trajectory length and mission construction time were slightly changed.
The total trajectory length generated by the system with A* path planning only was 76.84 km, which is significantly longer than the trajectory length from the human-generated flight paths, which was 62.95 km.
Trajectory length comparison and the error analysis are shown in figures below. Fig. 7 shows the comparison of trajectory lengths generated by the UAV-VLA system with path planning application and by an experienced operator.
Figure 7: The comparison of the trajectory lengths made by UAV-VLA with A* path planning algorithm and by an experienced operator.
The system performance was also evaluated the same as in the TSP only section V-A. The results are summarized and shown in Table II:
TABLE II: Comparison of RMSE Metrics for Different Methods on UAV-VLA with path planning
@@ V-B Results with the A* path planning algorithm
Mean 70.1541 262.675 403.305
Median 53.6504 267.013 366.284
Max 337.358 534.32 730.027
Error distribution of the system across different error metrics is drawn in Fig. 8.
Figure 8: The error of UAV-VLA with A* path planning algorithm in comparison with ground truth.
Flight comparison of the Human-made path and the path created by VLM and A* algorithm only is shown in the Fig. 9
@@
(b) UAV-VLA with Path planning
Figure 9: Comparison of flight plans generated by a human expert (a) and the UAV-VLA with Path planning only system (b).
V-C Results with TSP and A* Path Planning
The total trajectory length generated by the system with A* path planning and TSP (UAV-VLPA system) only was 51.27 km, which is shorter than the trajectory length from the human-generated flight paths, which was equal to 62.95 km. The system performance was also evaluated similarly to the TSP only and Path planning only sections. The results are summarized and shown in Table  III:
Trajectory length comparison and the error analysis are shown in figures below. Fig. 10 shows the comparison of trajectory lengths generated by the UAV-VLA system with path planning application and by an experienced operator.
Figure 10: The comparison of the trajectory lengths made by the UAV-VLPA system and by an experienced operator.
TABLE III: Comparison of RMSE Metrics for Different Methods on the UAV-VLPA system
@@ V-C Results with TSP and A* Path Planning
Mean 69.1918 223.936 348.24
Median 54.4354 170.953 291.855
Max 337.886 476.072 723.488
Error distribution of the system across different error metrics is drawn in Fig. 11. Example of the path created by the UAV-VLPA system is shown in Fig. 12.
Figure 11: The error distribution of the UAV-VLPA system in comparison with ground truth.
@@
(b) UAV-VLPA system
Figure 12: Comparison of flight plans generated by a human expert (a) and the UAV-VLPA system (b).
@@ VI Conclusion
Application of the Vision-Language Model and Traveling Salesman Problem Method significantly increases the speed of the large mission generation. Overall mission generation time for VLM was 11 times less and total trajectory length for TSP-Path planning was 24.4% shorter than in the human-made missions. 22/30 or 73.3% TSP-Path planning trajectories were shorter than the human-made ones. The Implemetation of the Traveling Salesman Problem resulted in a reduction of the trajectory length from 77.74 km to 51.34 km.
When applying the A* algorithm directly to raw data, the resulting path distance increases by 22%. This clearly demonstrates the inefficiency of using unprocessed data for path planning.
The integration of the TSP approach followed by A* path-planning resulted in an overall trajectory reduction of 18.5% compared to human-generated paths. Additionally, the system demonstrated superior performance, achieving the shortest path in 25 out of 30 test cases. This highlights the effectiveness of the combined approach in optimizing path planning and reducing travel distance. Furthermore, the system’s ability to consistently outperform human-generated paths underscores its potential for practical applications in autonomous navigation and mission planning.
@@ VII Future Work
The integration of global path planning through language requests represents a transformative tool for future applications in autonomous systems. By leveraging open satellite data alongside classical mathematical approaches, such as the Traveling Salesman Problem (TSP) and A* algorithm, we can unlock powerful opportunities for optimizing navigation in complex environments. Future work will focus on enhancing the system’s ability to process real-time satellite imagery and dynamically adapt to changing environmental conditions, further improving path accuracy and efficiency.

### v2

# UAV-VLPA*: Vision-Language Mission Planning with Global-Local Optimization for Scalable UAV Route Generation

@@ Abstract
We introduce UAV-VLPA* (Visual-Language-Path-and-Action), a vision-language-based system for autonomous UAV mission planning that generates flight paths directly from natural language instructions and satellite imagery. The system integrates semantic understanding with two key planning components: the Traveling Salesman Problem (TSP) for globally optimized waypoint ordering, and the A* algorithm for obstacle-aware local path refinement.
Tested on the diverse UAV-VLPA-nano-30 benchmark, UAV-VLPA* consistently outperforms human-designed trajectories. Our experiments show that TSP alone reduces the total trajectory length by over 33.9% when compared with the baseline, while the combined TSP + A* setup achieves a 51.27 km path, improving on human-generated flight routes by 18.5%. Error analysis across KNN, DTW, and sequential RMSE metrics further confirms the system’s reliability and precision.
Designed for scalable deployment, UAV-VLPA* supports offloading computationally intensive visual-linguistic processing to external servers, enabling lightweight onboard execution. The results demonstrate the system’s ability to autonomously plan efficient and safe UAV missions—without manual intervention—highlighting its potential for real-world applications in complex, large-scale environments.
Keywords: VLA; VLM; LLM-agents; VLM-agents; UAV; Navigation; Drone; Path Planning.
@@ I Introduction
Figure 1: The pipeline of the UAV-VLPA* system.
Unmanned aerial vehicles (UAVs) are increasingly used in diverse domains such as surveillance, agriculture, and disaster response [1]. As mission complexity grows, the need for intelligent and autonomous path planning becomes critical. Traditional approaches based on manual waypoint setup or pre-programmed routes [2], [3] are often inefficient, requiring human expertise and lacking scalability in complex and unknown environments.
Recent work has shown that integrating visual data and natural language commands with navigation modules can enable more autonomous and intuitive UAV control [4], [5], [6]. Our prior systems, UAV-VLA [7] and UAV-CodeAgents [8], translates natural language instructions into flight paths using satellite imagery leveraging zero-shot capabilities of powerful models without additional training. Building on this, we present UAV-VLPA*, an improved framework that integrates (1) semantic understanding via vision-language models, (2) global route optimization through the Traveling Salesman Problem (TSP), and (3) local obstacle avoidance using the A* algorithm.
UAV-VLPA* addresses key challenges in UAV navigation by generating efficient, shorter, and safer routes—without requiring manual flight data or expert-tuned instructions. The system autonomously constructs optimal flight plans directly from satellite imagery and user commands, significantly reducing planning time and human effort.
We evaluate our system using the UAV-VLPA-nano-30 benchmark [7], which features diverse real-world environments across the United States, including urban, suburban, rural, and natural scenes. It includes buildings, stadiums, lakes, roads, bridges, and more, captured in daylight during spring and summer. This diversity allows for robust evaluation of route planning capabilities in visually and structurally varied contexts.
Our contributions are as follows:
  • We propose UAV-VLPA*, combining Vision-Language understanding with TSP and A* for globally optimal and obstacle-aware UAV path planning.
  • We present a scalable instruction-driven navigation system that eliminates the need for manual flights or environment-specific tuning.
  • We demonstrate that UAV-VLPA* outperforms human experts and baselines in route efficiency on the UAV-VLPA-nano-30 benchmark.
@@ II Related Work
The introduction of Vision Transformers (ViT) [9],[10] marked a significant advancement in the development of full-fledged models capable of processing and integrating multiple types of input and output, including text, images, video, and more. Building on this progress, the leading companies introduced models like ChatGPT-4 Omni[11] and DeepSeek-VL [12], which can reason across audio, vision, and text in real time, enabling seamless multimodal interactions. To address the problem of objects finding in robotics applications, Allen Institute of AI introduced model Molmo, that can point the requested objects on an image[13].
The usage of the transformer-based models allowed the extensive developing of the new methods, benchmarks, and datasets for Vision Language Navigation and Vision Language Action tasks. Firstly, the problem of Aerial Visual Language Navigation was proposed by Liu et al. [14], where they introduced the Aerial VLN method together with AerialVLN dataset. In [6] Fan et al. described the simulator and VLDN system, that can support the dialog with an operator during the flight. Lee et al.[4] presented an extended dataset with geographical meta information (streets, squares, boulevards, etc.). The introduction of dataset was paired with the new approach for goal predictor. Zhang et al. [15] took a pioneering step by building a universal environment for embodied intelligence in an open city. The agents there can perform both VLA and VLN tasks together online. Gao et al. [16] presented a method, where a map was provided as a matrix to the LLM model. In that work was introduced the Semantic Topo Metric Representation (STMR) approach, that allowed to feed the matrix map representation into the Large Language Model. In [17] Wang et al. presented the benchmark and simulator dubbed OpenUAV platform, which provides realistic environments, flight simulation, and comprehensive algorithmic support. To address the problem of a flight support from the human point of view, Sautenkov et. al [2] suggested system to visually enhance the drone control with several video streams and object detection. The system helps the operator to conduct the surveillance tasks, however, doesn’t exclude him from the general pipeline.
@@ III System Overview
Originally, the vision language model (VLM) provided outputs by identifying point objects on an image in a sequential manner, from left to right and top to bottom, without considering obstacles or their spatial relationships [7]. To address these limitations, we propose the following steps.
@@ III-A System Workflow
    Figure 2: The buildings detected by VLM (purple colored dots)
@@
    Figure 3: The 2-Opt Local TSP Search path construction through points detected by VLM
@@
    Figure 4: The A* algorithm’s route generation considering the obstacle existence through optimal path directions
@@ III-C Path Planning and Obstacle Avoidance
Since A* algorithm is known as one of the most accurate [24] and fastest tools to create a path from one starting point to another goal point, it is chosen as the main algorithm for this work. To efficiently execute the A* algorithm in large-scale environments, the map is discretized into a uniform grid of equally sized cells, where each cell represents a node in the graph. The grid resolution is determined by balancing path fidelity—how accurately the path adheres to real-world terrain and obstacles—and computational efficiency—the number of nodes processed during the search. Considering the balance between smoothness of the trajectory and time-efficiency of the execution, the grid cell size is set to 5 x 5. (discretization step is 5 pixels).
@@ IV Experimental Setup
To evaluate the UAV-VLPA* system, we used the natural language instruction: “Create a flight plan for the quadcopter to fly around each building at a height of 100 m, return to home, and land at the take-off point. Avoid lakes.” The experiments were conducted on a high-performance workstation equipped with an NVIDIA RTX 4090 GPU (24GB VRAM) and an Intel Core i9-13900K processor. To accommodate hardware limitations, particularly memory constraints, the quantized 4-bit Molmo-7B-O BnB model [25] was utilized.
@@ V Experimental Results
V-A Evaluation of TSP Integration on UAV-VLA results
To assess the impact of incorporating the Traveling Salesman Problem (TSP) into UAV path optimization, we evaluated the system using the UAV-VLPA-nano-30 benchmark. TSP is a classical combinatorial optimization problem that seeks the shortest possible route visiting a set of locations exactly once and returning to the origin. In the context of UAV mission planning, solving TSP over high-level waypoints (e.g., buildings or mission targets) ensures a globally efficient route, minimizing energy consumption and flight time—critical for resource-constrained UAV operations.
The integration of TSP led to a significant reduction in total flight distance. The generated trajectory covered 51.34 km, compared to 63.89 km in the human-crafted routes reported in the UAV-VLA benchmark [7]. Notably, our implementation reduced the original trajectory (77.74 km) by over 33.9%, confirming the benefits of TSP-based waypoint reordering.
We further evaluated the system using RMSE-based metrics (the same as in the baseline [7]: KNN error, DTW error, and Sequential Interpolation error. Results are presented in Table I:
TABLE I: RMSE Metrics for UAV-VLA System with TSP Integration
@@ V-A Evaluation of TSP Integration on UAV-VLA results
Mean 45.16 259.07 354.46
Median 26.89 235.48 308.91
Max 336.22 719.48 734.32
Figures 5 and 6 visualize the comparison in trajectory length and error metrics, respectively. TSP-enhanced flight plans consistently outperformed human-designed ones in terms of efficiency while maintaining competitive accuracy across trajectory alignment metrics.
Figure 5: Trajectory length comparison: UAV-VLA with TSP vs. human expert.
Figure 6: Error distribution of UAV-VLA with TSP against ground truth.
V-B Evaluation of A* Path Planning
In this and subsequent experiments, the human-generated trajectory was updated to reflect the new constraint of avoiding lakes, which slightly altered both the total trajectory length and mission construction time.
Using only the A* algorithm for low-level path generation—without TSP for waypoint reordering—the system produced a trajectory of 76.84 km. This is notably longer than the updated human-planned route of 62.95 km. The increase suggests that while A* is effective at generating obstacle-avoiding paths, its sequential application without global optimization (e.g., TSP) leads to inefficiencies in overall mission planning.
Figure 7 shows the comparative trajectory lengths, and Table II presents the corresponding RMSE error metrics.
Figure 7: Trajectory length comparison: UAV-VLA with A* vs. human expert.
TABLE II: RMSE Metrics for UAV-VLA System with A* Path Planning
@@ V-B Evaluation of A* Path Planning
Mean 70.15 262.68 403.31
Median 53.65 267.01 366.28
Max 337.36 534.32 730.03
Figure 8: Error distribution of UAV-VLA with A* path planning.
@@
(b) UAV-VLA with A*
Figure 9: Flight plans comparison: (a) human expert vs. (b) UAV-VLA with A* path planning.
V-C Combined TSP and A* Path Planning: UAV-VLPA* System
When combining both TSP (for waypoint reordering) and A* (for fine-grained obstacle avoidance), the UAV-VLPA* system achieved the best performance. The resulting trajectory was 51.27 km, outperforming the updated human reference trajectory of 62.95 km by 18.5%.
This result highlights the complementary strengths of the two algorithms: TSP provides a globally optimal waypoint visitation order, while A* ensures feasible and safe traversal between those waypoints.
Figure 10: Trajectory length comparison: UAV-VLPA* system vs. human expert.
TABLE III: RMSE Metrics for UAV-VLPA* System (TSP + A*)
@@ V-C Combined TSP and A* Path Planning: UAV-VLPA* System
Mean 69.19 223.94 348.24
Median 54.44 170.95 291.86
Max 337.89 476.07 723.49
Figure 11: Error distribution of UAV-VLPA* system with TSP and A* integration.
@@
(b) UAV-VLPA*
Figure 12: Flight plans comparison: (a) human expert vs. (b) the UAV-VLPA* system (TSP + A*).
In all of the given approaches, the total execution time of the whole pipeline took on HPC no more than 3 minutes, whereas human-made trajectory constructions took in average 35 minutes.
@@ VI Conclusion
In this paper, we introduced UAV-VLPA*, a vision-language-based UAV mission planning framework that integrates natural language understanding, satellite imagery, TSP-based global optimization, and A* obstacle-aware path planning. Our goal was to generate efficient, safe, and scalable UAV flight plans directly from user instructions—without requiring manual flight demonstrations or labor-intensive path design.
We evaluated our system using the UAV-VLPA-nano-30 benchmark, which features a diverse set of environments across the United States. Through extensive experiments, we demonstrated the effectiveness of each component (TSP, A*, and their combination) in improving route efficiency and mission execution quality.
Key findings of this research are:
  • TSP Integration on the results of the baseline UAV-VLA [7]: Reduced trajectory length by 33.9% compared to the original baseline, outperforming both human-made and previous system-generated routes.
  • A* Path Planning: Enabled obstacle avoidance in local path segments but resulted in longer global paths without TSP, showing the importance of combining global and local planning.
  • UAV-VLPA* (TSP + A*): Achieved the best performance with a trajectory length of 51.27 km, improving upon the human reference path by 18.5%, while maintaining low error across multiple trajectory alignment metrics.
  • Scalability and Generalization: The system successfully handled varied geographic regions and structural layouts using open satellite imagery and flexible language input, with no environment-specific tuning.
Overall, UAV-VLPA* represents a step forward in intelligent UAV mission generation, offering a practical and scalable solution for natural language-driven flight planning in complex real-world scenarios. Future work will explore integrating 3D terrain data, dynamic obstacle handling, and real-time adaptation to further enhance mission robustness.
@@ VII Future Work
The integration of natural language-driven global path planning marks a promising direction for advancing autonomous aerial systems. By combining open-access satellite imagery with classical optimization methods like the Traveling Salesman Problem (TSP) and A* algorithm, our current system lays the groundwork for efficient mission generation in complex and large-scale environments.
Future improvements will focus on addressing real-time operational challenges. Specifically, we aim to:
  • Enhance computational efficiency for real-time deployment by optimizing the system’s inference pipeline and reducing latency in path generation.
  • Support dynamic environmental adaptation by integrating up-to-date satellite or UAV-captured visual data, enabling reactive re-planning in response to evolving conditions such as temporary obstacles or weather changes.
  • Deploy real-world flight experiments where onboard UAV systems will communicate with a remote server. In this setup, high-compute modules (Vision-Language Model and Large Language Model) will be offloaded to the server, while onboard components will handle navigation and control, ensuring lightweight execution and real-time responsiveness.
  • Expand to 3D environments, incorporating terrain elevation data and urban structure heights to enable more realistic and safe UAV navigation in dense or uneven terrains.
By addressing these aspects, future iterations of UAV-VLPA* will be better positioned for real-world deployment, with improved robustness, scalability, and mission execution in dynamic, large-scale aerial scenarios.